OpenStudy (anonymous):

if r=cube root of A^2B/C then log r can be represented by

OpenStudy (kamille):

\[r=\sqrt[3]{\frac{ A ^{2B} }{ C}}\] is it your problem?

OpenStudy (anonymous):

the b isnt part of the exponent

OpenStudy (anonymous):

\[r=\sqrt[3]{A ^{2}B/C}\]

OpenStudy (anonymous):

thats what it looks like. well as close as i could get it :)

OpenStudy (anonymous):

\[\log r = \log (A ^{2}B/C)^{1/3}\]

OpenStudy (anonymous):

can you explain how you got that?

OpenStudy (anonymous):

If r=cube root of A^2B/C Then log r = log cube root of A^2B/C and since \[\sqrt[3]{x} = x ^{1/3}\] then, log r = log (A^2B/C)^1/3

OpenStudy (anonymous):

which can also be expressed as log r = 1/3 [ log (A^2B/C) ]

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

ur welcome :)